### v2

# Towards Fine-Grained Citation Evaluation in Generated Text: A Comparative Analysis of Faithfulness Metrics

Abstract
Large language models (LLMs) often produce unsupported or unverifiable content, known as “hallucinations.” To mitigate this, retrieval-augmented LLMs incorporate citations, grounding the content in verifiable sources. Despite such developments, manually assessing how well a citation supports the associated statement remains a major challenge. Previous studies use faithfulness metrics to estimate citation support automatically but are limited to binary classification, overlooking fine-grained citation support in practical scenarios. To investigate the effectiveness of faithfulness metrics in fine-grained scenarios, we propose a comparative evaluation framework that assesses the metric effectiveness in distinguishing citations between three-category support levels: full, partial, and no support. Our framework employs correlation analysis, classification evaluation, and retrieval evaluation to measure the alignment between metric scores and human judgments comprehensively. Our results show no single metric consistently excels across all evaluations, revealing the complexity of assessing fine-grained support. Based on the findings, we provide practical recommendations for developing more effective metrics.
Towards Fine-Grained Citation Evaluation in Generated Text:
A Comparative Analysis of Faithfulness Metrics
Weijia Zhang1  Mohammad Aliannejadi1  Yifei Yuan2  Jiahuan Pei3 Jia-Hong Huang1  Evangelos Kanoulas1 1University of Amsterdam  2University of Copenhagen 3Centrum Wiskunde & Informatica w.zhang2@uva.nl
Figure 1: An example of partial support in citation evaluation. Inconsistent metric scores are observed when assessing the statement with three faithfulness metrics.
1 Introduction
Large language models (LLMs) often generate content known as “hallucinations” Li et al. (2022); Ji et al. (2022); Zhang et al. (2023b), which contradicts established knowledge or lacks verification from reliable sources. Mainstream studies Bohnet et al. (2022); Gao et al. (2023a) aim to mitigate this by using retrieval-augmented LLMs to generate responses with in-line citations that provide supporting evidence. One primary challenge is to assess how well a citation supports its statement, as manual evaluation is labor-intensive and time-consuming. Automated citation evaluation has been explored to reduce reliance on human assessments Gao et al. (2023b); Li et al. (2024b). To this end, faithfulness evaluation metrics are employed as proxies to automatically estimate the citation support Xia et al. (2024); Li et al. (2024a). These metrics measure the faithfulness between model-generated and sourced text, which aligns closely with the objectives of automated citation evaluation.
Prior studies in faithfulness metrics have primarily limited this task to a binary classification problem Tahaei et al. (2024); Huang et al. (2024d), where faithfulness metrics are leveraged to determine whether a citation supports the associated statement. However, this binary approach fails to capture the fine-grained citation support encountered in real-world applications. For instance, in Figure 1, a retrieval-augmented LLM generates a response with multiple citations given a query. A human assessor labels the first citation as “partial support” since it only supports “the most humid place in Australia is Macquarie Island” but not “which is located in the Southern Ocean off the coast of Tasmania.” This partial support scenario causes noticeable inconsistencies across three different faithfulness metrics. Therefore, there is a significant research need to evaluate the effectiveness of faithfulness metrics in accurately distinguishing citations in such fine-grained support scenarios.
To address this issue, we propose a comparative evaluation framework for assessing the metric effectiveness in fine-grained support scenarios. In our framework, we define “support levels” as the extent to which a citation supports the associated statement Liu et al. (2023); Yue et al. (2023). Specifically, we consider a three-category support level scenario: full, partial, and no support. These categories indicate whether a citation fully, partially or does not support the associated statement, respectively. To comprehensively assess the metric effectiveness, we measure the alignment between metric scores and human judgments with three types of evaluation protocols: 1) Correlation analysis:we employ it to measure how well metric scores correlate with human judgments. 2) Classification evaluation:we conduct a classification evaluation to assess the metrics’ capability to distinguish citations based on their support levels. 3) Retrieval evaluation:we undertake a retrieval evaluation to assess the metric effectiveness in ranking citations according to their support levels. This is motivated by the observation that the previous two evaluation protocols assume citations are within statements, which is not always valid in practice Asai et al. (2024). In such cases, faithfulness metrics are adapted to perform post-hoc retrieval, aiming to retrieve potential citations from a candidate pool Kang et al. (2023); Gou et al. (2024). Thus, retrieval evaluation is crucial for determining the practical utility of these metric adaptations.
In our experiments, we assess various widely used faithfulness metrics, categorizing them into similarity-based, entailment-based, and LLM-based metrics. We find that: 1) No single faithfulness metric consistently outperforms others across three evaluation protocols, suggesting that these protocols are complementary and should be integrated for a comprehensive evaluation of metric performance; 2) The best-performing metrics show promise in distinguishing some support scenarios but struggle with others. This highlights the inherent complexities of automated citation evaluation. 3) Similarity-based metrics surpass best-performing entailment-based metrics in retrieval evaluation. This indicates that entailment-based metrics exhibit higher sensitivity to noisy data, which is introduced by irrelevant documents in such scenarios.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to systematically investigate the effect of fine-grained support levels on faithfulness metrics in the task of automated citation evaluation.
  • We propose a comparative evaluation framework to assess the alignment between metric scores and human judgments. This framework includes three evaluation protocols to comprehensively evaluate the metric performance.
  • Our experimental results demonstrate the best-performing faithfulness metrics still struggle to identify partially supporting citations, underscoring the inherent challenges of automated citation evaluation. Based on our findings, we offer practical recommendations for the development of more effective metrics.
Figure 2: The overview of the proposed comparative evaluation framework. A faithfulness metric assigns scores to given statements and their corresponding citations. Subsequently, our framework comprehensively assesses the alignment between these metric scores and human judgments by employing correlation analysis, classification, and retrieval evaluation.
2 Related Work
Faithfulness Evaluation Metrics
Faithfulness evaluation metrics are crucial for assessing the factual consistency of text generated by models relative to the source text. It receives great interest within the field of natural language generation (NLG) Huang et al. (2019, 2021b); Zhang et al. (2021); Huang et al. (2022, 2023); Zhang et al. (2023a); Huang et al. (2024b, c); Zhu et al. (2024), particularly in abstractive summarization Maynez et al. (2020); Kryscinski et al. (2020); Huang and Worring (2020); Huang et al. (2021a); Zhang et al. (2024b, c). In general, faithfulness metrics are categorized into three types: entailment-based, similarity-based, and QA-based metrics. Entailment-based metrics employ natural language inference (NLI) models to determine if the source text entails the generated text Falke et al. (2019); Laban et al. (2022); Honovich et al. (2022); Zha et al. (2023). Similarity-based metrics, such as BERTScore Zhang et al. (2020) and BARTScore Yuan et al. (2021), quantify text similarity and have demonstrated robust performance in faithfulness evaluation Pagnoni et al. (2021); Honovich et al. (2022). QA-based metrics utilize a combination of question generation and question answering to estimate faithfulness levels Durmus et al. (2020); Wang et al. (2020); Scialom et al. (2021); Fabbri et al. (2022). In this work, we exclude QA-based metrics from our work, following recent works suggesting the challenging limitations in these metrics Kamoi et al. (2023). We focus on the extrinsic evaluation of faithfulness metrics against human judgments in scenarios requiring fine-grained citation support.
Citation Evaluation
Citation evaluation seeks to enhance the trustworthiness of retrieval-augmented LLMs by verifying the support provided by citations to the generated statements Rashkin et al. (2023); Yue et al. (2023); Huang and Chang (2023); Huang et al. (2024a); Zhang et al. (2024a). Given the labor-intensive nature of manual citation evaluation, there has been a shift towards automated approaches to reduce dependence on human evaluation. Since the goals of automated citation evaluation align closely with faithfulness evaluation in NLG, faithfulness metrics are employed to verify whether a citation supports the corresponding statement Li et al. (2024c); Sun et al. (2023); Ye et al. (2024); Li et al. (2024d); Shen et al. (2024); Huang et al. (2024d). Despite their widespread usage, the effectiveness of these metrics in more practical fine-grained citation support scenarios, such as those involving partial support by citations, has not been adequately addressed. Questions remain about the metrics’ capability to differentiate citations in these fine-grained scenarios. This work addresses these gaps by examining the effectiveness of faithfulness metrics across three distinct levels of citation support: full, partial, and no support.
3 Evaluation Framework
In this section, we introduce the proposed comparative evaluation framework. We begin by formalizing the task of automated citation evaluation. Subsequently, we detail three distinct evaluation protocols within this framework, ensuring a comprehensive assessment in alignment between faithfulness metrics and human judgments. Our framework is demonstrated in Figure 2.
3.1 Task Formulation
The objective of automated citation evaluation is to automatically quantify the support level of a citation based on the citation and its associated statement. In this work, we assume access to a dataset for automated citation evaluation, comprising pairs of statements and their corresponding citations, denoted as $(s_{i},c_{i})$. Each $s_{i}$ is a statement from the set $S$ of all statements produced by an LLM and each $c_{i}$ is a citation from a set $C$ of citations returned by the LLM. We categorize the citations into three distinct support levels: full, partial, and no support. We adopt the definition of these support levels from Liu et al. (2023):
  • Full Support (FS): the citation fully supports every detail in the statement.
  • Partial Support (PS): the citation supports certain aspects of the statement, while other details remain unsupported or are contradicted.
  • No Support (NS): none of the content in the statement is supported by the citation. For instance, the citation is entirely irrelevant or contradicts the statement.
To this end, without loss of generality, we define a faithfulness metric as a scoring function, denoted as $F(s_{i},c_{i})\rightarrow R^{+}$. For any given statement $s_{i}$ and its associated citation $c_{i}$, this scoring function provides a numeric score that indicates the extent of support provided by the citation to the statement.
3.2 Evaluation Protocols
The objective of evaluation protocols is to comprehensively assess the extent to which metric scores align with human judgments. In this work, we assess this alignment across three distinct dimensions: correlation, classification performance, and retrieval effectiveness.
3.2.1 Correlation Analysis
[…]
3.2.2 Classification Evaluation
In addition to correlation analysis, we perform classification evaluation to determine the metric effectiveness in discriminating citations based on their support level. Specifically, the metrics need to categorize a citation into one of three support levels: FS, PS, NS. Notably, existing faithfulness metrics do not apply to this three-way classification scenario, as they are unable to accurately determine the extent to which a statement is partially supported by its corresponding citation Laban et al. (2022). To address this issue, we adopt a one-vs-one strategy, by effectively decomposing the three-way classification into three binary classification task settings: (i) Full Support vs. No Support (FS-vs-NS), (ii) Full Support vs. Partial Support (FS-vs-PS), and (iii) Partial Support vs. No Support (PS-vs-NS). For each binary classification task setting, we construct a specialized dataset comprising only instances with the corresponding binary support levels derived from the original dataset. We assess the performance of metrics on these tailored binary datasets using standard binary classification evaluation metrics. The overall metric performance is then computed by averaging the results across all binary tasks.
3.2.3 Retrieval Evaluation
[…]
4 Faithfulness Metrics
In our experiments, we evaluate diverse faithfulness evaluation metrics, dividing them into similarity-based, entailment-based, and LLM-based metrics. Similarity-based metrics assess the support levels mainly based on the degree of similarity between the citation and the associated statement. Entailment-based metrics leverage pre-trained NLI models to estimate the support levels. LLM-based metrics directly prompt LLMs to measure the support levels.
4.1 Similarity-Based Metrics
BERTScore Zhang et al. (2020) adopts BERT Devlin et al. (2019) to measure semantic similarity between a pair of text by aggregating cosine similarity among token-level BERT representation without further fine-tuning. We report the precision version of BERTScore since it correlates more with human judgments in faithfulness evaluation Pagnoni et al. (2021), We use recommended deberta-xlarge-mnli He et al. (2021) as the backbone model.
BARTScore Yuan et al. (2021) adopts BART Lewis et al. (2020) to measure the similarity between two texts based on conditional log-likelihood of generating target text from source text. In our experiments, we leverage the faithfulness version of BARTScore, in which we treat the citation and the statement as the source and target text, respectively. We use the BART model fine-tuned on the CNN/DailyMail dataset Hermann et al. (2015) as the backbone model.
4.2 Entailment-Based Metrics
FactCC Kryscinski et al. (2020) is a BERT-based model to verify whether a generated text is faithful to a source text, which is fine-tuned on synthetic training data containing simulated examples with different factual errors Kryscinski et al. (2020).
SummaC  Laban et al. (2022) is a RoBERTa-based model Liu et al. (2019) fine-tuned on NLI datasets. This metric splits source and generated texts into sentences, computes entailment scores for each pair, and aggregates these scores to obtain the final faithfulness score. It has two variants: (i) SummaCZS is a zero-shot version that is only pre-trained on NLI datasets; (ii) SummaCConv adds extra convolutional layers and is further fine-tuned on synthetic training data proposed in Kryscinski et al. (2020).
AutoAIS Honovich et al. (2022) is a T5-11B Raffel et al. (2020) model trained on a collection of NLI datasets, which is commonly used in recent automated citation evaluation. As the original output of AutoAIS is a numeric, either “1” (faithful) or “0” (unfaithful), we utilize the generated token probability of “1” as the predicted metric score.
AlignScore Zha et al. (2023) further fine-tunes a RoBERTa-based model Liu et al. (2019) with a unified alignment loss function. To this end, a unified dataset containing a variety of related natural language processing datasets has been collected. In this work, we adapt the large version as it demonstrates the best performance.
4.3 LLM-Based Metrics
In addition to established faithfulness metrics, we utilize LLMs as faithfulness evaluators for comparison. Specifically, we introduce two prompting methods as follows: (i) Discrete scoringprompts the LLM to assign discrete scores from the set ${0,1,2}$ for a given statement and its citation, where 00, $1$, and $2$ indicate no support, partial support, and full support, respectively; (ii) Continuous scoringprompts the LLM to assign continuous scores in the range $[0,1]$ for a given statement and its citation. Here, $1$ indicates full support, 00 indicates no support, and values between 0 and 1 indicate partial support.
In the experiments, we employ the latest version of GPT-3.5 (gpt-3.5-turbo-0125) as the base model. Moreover, we utilize the chain of thought (CoT) method Wei et al. (2022); Kojima et al. (2022) to enhance the reasoning capabilities of the LLM. We use GPT-3.5-DIS and GPT-3.5-CON to denote GPT-3.5 using discrete and continuous scoring methods, respectively. The detailed prompts are shown in Appendix A.
Human Judgment # Statement-Citation Pair
Full Support 6,616
Partial Support 1,445
No Support 4,620
Total 12,681
Table 1: Data statistics of the GenSearch dataset. Each pair has been annotated by human assessors based on three categories: full, partial, and no support.
5 Experiments
In this section, we describe the dataset used in the experiments. Subsequently, we discuss the evaluation metrics incorporated within our proposed framework, which assess the performance of faithfulness metrics in alignment with human judgments.
5.1 Datasets
In our experiments, we utilize the GenSearch dataset Liu et al. (2023) as our evaluation benchmark, which consists of data from generative search engines (GSE) like BingChat. [^1] https://www.bing.com/chat These GSEs represent commercial applications of retrieval-augmented LLMs. As depicted in Figure 1, each example includes a user query and a corresponding response generated by the GSE. The user queries are sourced from various QA datasets Fan et al. (2019); Kwiatkowski et al. (2019). Each response consists of multiple statements, each containing in-line citations linking to web documents. Notably, these statements are supported by one or more citations. For this benchmark, human assessors are enrolled to annotate each statement-citation pair based on the degree to which the citation supports the associated statement.
Data Statistics
The GenSearch dataset comprises a total of $12,681$ statement-citation pairs. For each pair, human assessors categorize the citation into one of three categories of support levels: full, partial, or no support. The details of data statistics are shown in Table 1. Notably, for citations classified under the full or partial support categories, human assessors additionally extract explicit evidence sentences from the citation that support the associated statement.
Data Processing
While the GenSearch dataset aligns well with our research objectives, we encounter a significant challenge: the extensive length of most citations within the dataset. These citations often comprise a web document with thousands of words, far exceeding the maximum input capacity of most faithfulness metrics, which is limited to $512$ tokens. This limitation necessitates input truncation, potentially compromising the reliability of faithfulness metrics. To mitigate this issue, we adopt a strategy similar to previous studies Zha et al. (2023). Specifically, we segment each cited document into shorter text chunks, with a maximum length of $150$ words per chunk. These text chunks, along with their corresponding statements, serve as the inputs for faithfulness metrics to predicted metric scores. Furthermore, to determine human judgments for the text chunks, we employ the Jaccard similarity index to identify text chunks containing human-annotated evidence sentences, classifying them as either fully or partially supporting text chunks.
5.2 Evaluation Metrics
We report Pearson, Spearman, and Kendall coefficients for correlation analysis, as recommended by previous research Pagnoni et al. (2021). In terms of classification evaluation, following previous studies Honovich et al. (2022); Ma et al. (2023), we report the macro-averaged Receiver Operating Characteristic-Area Under Curve (ROC-AUC) score, as it obviates the need for manual threshold setting for each binary classification task. For retrieval evaluation, we report standard normalized discounted cumulative gain (NDCG@n) scores where $n\in\{5,10,20\}$.
Metric Pearson Spearman Kendall
LLM-based
   GPT-3.5-CON 0.023 0.057 0.035
   GPT-3.5-DIS 0.101 0.181 0.128
Entailment-based
   FactCC 0.121 0.199 0.140
   SummaCZS 0.364 0.180 0.137
   SummaCConv 0.565 0.444 0.342
   AlignScore 0.585 0.488 0.393
   AutoAIS 0.638 0.639 0.547
Similarity-based
   BERTScore 0.542 0.227 0.170
   BARTScore 0.598 0.235 0.176
Table 2: Correlation coefficients between human-annotated support levels and metric scores on the GenSearch dataset. The best and second-best correlations are marked in bold and underline, respectively.
Category Metric FS-vs-NS FS-vs-PS PS-vs-NS Overall
LLM-based GPT-3.5-CON 54.80 54.13 51.60 53.51
GPT-3.5-DIS 57.84 52.79 55.48 55.37
[…]
SummaCZS 78.60 72.96 58.67 70.08
SummaCConv 85.01 78.74 61.84 75.20
AlignScore 90.79 81.41 69.78 80.66
AutoAIS 92.61 82.31 73.90 82.94
[…]
BERTScore 91.55 75.94 78.72 82.07
Table 3: Classification performance of faithfulness metrics regarding ROC-AUC score (%) on the GenSearch dataset. The overall performance is the macro-averaged performance of three binary classification settings. The best and second-best scores are marked in bold and underline, respectively.
6 Results and Analyses
In this section, we discuss the performance of faithfulness metrics across three distinct evaluation protocols. Subsequently, we conduct a qualitative analysis through case studies.
6.1 Correlation Results
The correlation results are demonstrated in Table 2. The following observations can be made: 1) The best-performing metrics reveal moderate correlations when analyzed using the Pearson coefficient. For instance, AutoAIS achieves the highest Pearson coefficient, recording a value of $0.638$, largely surpassing the second-best BARTScore, which posts a coefficient of $0.598$. 2) There is notable variation in correlation trends among high-performing metrics. BARTScore shows the second-best Pearson correlation but much lower Spearman and Kendall correlations. This divergence likely arises from the Pearson coefficient’s assumption of linear relationships between two variables, which is often invalid in automated citation evaluation. 3) Similarity-based metrics generally show lower Spearman and Kendall correlations compared to Pearson. For instance, BERTScore has a substantial Pearson correlation of $0.542$ but lower Spearman and Kendall correlations of $0.227$ and $0.170$. This indicates that similarity-based metrics do not align well with human judgments, highlighting their limitations in fine-grained support scenarios. 4) LLM-based metrics show little correlation with human judgments among all correlation coefficients, with the correlation of the GPT-3.5-CON metric being almost zero. This finding suggests a negligible relationship between LLM-based metric scores and human judgments. Furthermore, the GPT-3.5-DIS metric significantly outperforms GPT-3.5-CON, highlighting that more fine-grained support levels present greater challenges in correlation analysis.
Figure 3: Retrieval performance of faithfulness metrics regarding NDCG@n scores on the GenSearch dataset. Note that we assign relevance labels $2$, $1$, and 00 to full, partial, and no support, respectively (shown in the color).
6.2 Classification Results
Table 3 presents the results of the classification evaluation. The observations can be summarized as follows: 1) Among all three binary classification task settings, most faithfulness metrics demonstrate superior performance in the FS-vs-NS setting. Notably, entailment-based AutoAIS achieves the highest ROC-AUC score of $92.61$, which shows significant discriminability between full support and no support instances. This can be attributed to its much more extensive parameters compared to other entailment-based metrics. 2) We observe the performance decline across the other two settings (i.e. FS-vs-PS and PS-vs-NS). For instance, when comparing the FS-vs-NS and PS-vs-NS settings, the ROC-AUC score of AutoAIS diminishes from $92.61$ to $73.90$. This decline indicates that even the best-performing metric struggles with granular sensitivity to varying levels of support. 3) While entailment-based AutoAIS generally surpasses other metrics, it is outperformed by similarity-based BERTScore in the PS-vs-NS setting. Interestingly, while most metrics perform worst in this setting, BERTScore shows its least effectiveness in FS-vs-PS. This highlights the unique prediction behaviors of different metrics across binary classification settings. 4) The performance of LLM-based metrics significantly lags behind other metrics. For instance, GPT-3.5-DIS achieves only a ROC-AUC score of $57.84$ in the FS-vs-NS setting, markedly lower than the best-performing AutoAIS, which achieves a ROC-AUC score of $92.61$. Furthermore, the overall performance of LLM-based metrics approaches random guessing. This underscores the inefficacy of LLM-based metrics in distinguishing fine-grained support levels.
6.3 Retrieval Results
Figure 3 presents the results of the retrieval evaluation. The key findings are as follows: 1) Similarity-based metrics, BARTScore and BERTScore, outperform other entailment-based metrics in all NDCG@n scores. For instance, entailment-based AutoAIS exhibits weaker NDCG@5 scores than BARTScore. This is likely because entailment-based metrics are more sensitive to noisy information than similarity-based metrics, as many irrelevant documents exist in retrieval scenarios. It suggests the need for the robustness improvements of metrics in post-hoc retrieval scenarios. 2) The best-performing BERTScore achieves more than twice the NDCG@n scores compared to LLM-based metrics. This result suggests that LLM-based metrics are ineffective in ranking documents with higher support levels. A plausible explanation is that LLM-based metrics lack fine-grained sensitivity to variations in support levels. Interestingly, our observations reveal that GPT-3.5-CON surpasses GPT-3.5-DIS, highlighting the advantage of fine-grained scoring methods in retrieval evaluation. 3) NDCG@n scores effectively capture the performance variations as the number of text chunks increases. For instance, as the chunk count increases, BARTScore shows a marginal performance improvement, while FactCC exhibits a more pronounced enhancement.
Error Reason Example
The citation does not explicitly mention coreference. Statement: Others believe that performance-enhancing drugs should be allowed in sports. Citation: However, if children are allowed to train as professional athletes, then they should be allowed to take the same drugs, provided that they are no more dangerous than their training is … Human Judgment: full support Metric Score: 0.055 (no support)
The complex statement includes independent claims. Statement: Love leads to growth while being in love is about ownership Citation:Growing to love the real person and accepting who they are, with both strengths and weaknesses, can make a wonderful difference in your relationship,” McCoy says … Human Judgment: partial support Metric Score: 0.0004 (no support)
The citation is semantically similar but non-supporting. Statement: Carpal tunnel syndrome can be treated with various methods, including wrist splinting, anti-inflammatory medication, and surgery. Citation: If diagnosed and treated early, the symptoms of carpal tunnel syndrome can often be relieved without surgery. If your diagnosis is uncertain or if your symptoms are mild, your doctor will recommend nonsurgical treatment first … Human Judgment: no support Metric Score: 0.52 (partial support)
Table 4: Case study of the faithfulness metric AutoAIS. Green phrases indicate supported content in the statement and corresponding supporting evidence. Red phrases indicate unsupported content in the statement and corresponding misleading information in the citation.
6.4 Case Study
Table 4 presents three cases of AutoAIS. In the first example, where human judgment indicates full support. AutoAIS incorrectly assigns a very low score. This may be due to the lack of explicit mention of drug coreference in the cited text chunk. This indicates coreference resolution remains a significant challenge in automated citation evaluation. In the second example, where human judgment indicates partial support. The complex statement implicitly contains two independent claims that require verification. However, the provided citation fails to offer sufficient evidence, resulting in an almost zero metric score. In the third example, where human judgment indicates no support. The given citation is semantically similar to the statement, leading to a metric score of partial support. Despite this semantic similarity, specific treatments mentioned in the statement, such as wrist splinting, are not explicitly referenced in the citation.
7 Discussions
Overall, our results across three evaluation protocols indicate that the evaluation protocols are complementary and should be integrated for a comprehensive assessment of metrics. Based on the evaluation results, we further propose the following practical recommendations to develop more effective metrics for automated citation evaluation: 1) Development of training resources:motivated by the observation that the best-performing metrics still struggle with identifying partial support, we recommend the development of training resources that include fine-grained support level annotations. These resources could significantly enhance the metrics’ fine-grained sensitivity to varying support levels; 2) Introduction of contrastive learning:to improve the robustness of metrics in post-hoc retrieval scenarios, we recommend fine-tuning metrics using contrastive learning frameworks. This method has demonstrated effectiveness across various information retrieval tasks Izacard et al. (2022). 3) Development of more explainable metrics:traditional faithfulness metrics often only provide final scores without sufficient explainability Xu et al. (2023). This limitation hinders a deeper understanding of the models’ behavior. Therefore, it is crucial to develop more explainable faithfulness metrics, potentially using large language models (LLMs).
8 Conclusion
We propose a comparative evaluation framework to explore the efficacy of faithfulness metrics beyond the binary scenario by examining three levels of citation support. Our framework employs correlation analysis, classification evaluation, and retrieval evaluation to measure the alignment between metric scores and human judgments. Experimental results reveal that no single metric consistently excels across all evaluation protocols, indicating the complexity of automated citation evaluation and the limitations of existing faithfulness metrics. We provide practical suggestions based on the findings.
Limitations
In this work, we consider a citation that explicitly contains human-annotated evidence as the fully supporting citation for each statement. However, for some complex statements, their evidence is distributed among multiple citations. For instance, about $2\%$ statements on the GenSearch dataset require multiple citations to be fully supported. Also, we focus on statement-level citation evaluation. Since answer-level citation evaluation is much more complicated and requires proper aggregation methods, we leave this exploration as future work. We do not evaluate QA-based faithfulness metrics as a recent study shows that such metrics have some fundamental issues, such as failing to localize errors Kamoi et al. (2023). However different findings could be explored with QA-based metrics. Finally, there is another line of work, in which they break down a complex statement into multiple atomic facts.
Ethical Considerations
We realized there are some risks in exploring citation evaluation for LLM-generated text. Since we have used publicly available datasets and open-source implementation of faithfulness metrics, we carefully avoid potential ethical problems caused by datasets or open-source codes. As we address the issue of the effectiveness of faithfulness metrics for LLM-generated text, concerning hallucination. We acknowledged the hallucinated text generated by LLMs may contain potentially harmful or misleading information. Our final goal is to mitigate such hallucination issues, which should support the discussion around hallucinations of LLMs and all ethical aspects around them.
Acknowledgments
This research was supported by the China Scholarship Council (CSC) under grant number 202008440470. The views expressed in the content belong solely to the authors and may not reflect the perspectives or endorsements of their respective employers or sponsors.
Appendix A Details of Prompts
Details of GPT-3.5 prompts used in the experiments are shown in Table 5.
Prompt Name Prompt Content
Discrete Scoring Instruction: Your task is to quantify how well a provided citation supports a given statement. You should predict a discrete score from the set $\{0,1,2\}$, where 00, $1$, $2$ represent that the statement is not supported, partially supported, and fully supported, respectively. Let’s think step by step. Statement: {statement} Citation: {cited text chunk} Prediction:
Continuous Scoring Instruction: Your task is to quantify how well a provided citation supports a given statement. You should predict a continuous score between 0 and 1 (inclusive), where 0 is not supported, 1 is fully supported, and a float value between 0 and 1 is partially supported. Let’s think step by step. Statement: {statement} Citation: {cited text chunk} Prediction:
Table 5: Detailed prompts for discrete and continuous scoring methods.